OpenStudy (anonymous):

solve for x 20x^2-7x=6

OpenStudy (solomonzelman):

\[20x^2-7x=6\]\[20x^2-7x\color{red} { -6 }=6\color{red} { -6 }\]\[20x^2-7x-6=0\]Now use the quadratic formula.

OpenStudy (anonymous):

x=−2/5,3/4

OpenStudy (solomonzelman):

\[\huge\color{green}{\frac{-b±\sqrt{b^2-4ac}}{2a}}\]

OpenStudy (solomonzelman):

We are NOT giving out plain answers!!!!! @sirizarry78!!!

OpenStudy (anonymous):

x=−2/5,3/4

OpenStudy (solomonzelman):

Guys stop giving out answers, the website is not for that!

OpenStudy (solomonzelman):

Do you understand how to do it, @mili85?

OpenStudy (solomonzelman):

(sorry, @milli85

OpenStudy (anonymous):

nt really

OpenStudy (anonymous):

im lost wen it comes to math

OpenStudy (solomonzelman):

\[\huge\color{red}{a}x^2+\huge\color{red}{b}x+\huge\color{red}{c}=0\]\[COMPARE~~~YOUR'S~~~TO~~THIS.\]\[\huge\color{red}{20}x^2\huge\color{red}{-7}x\huge\color{red}{-6}=0\] see the a b c? PLUG them into the quadratic formula.